OpenStudy (anonymous):

Can someone check this? Part 1 1. Solve. y + 12 = 7 –5<< –4 4 5 Question 2.2. Solve. 46 – 2x = 8 19<< –19 27 –27 Question 3.3. Solve. 6x = –42 –9 –8 –7<< 9 Question 4.4. What number must multiply each side of the equation 2/3x=16 to produce the equivalent equation x=24 3/2 2/3 << -2/3 -3/2 Question 5.5. Solve. -1/3x=12 -36<< -4 -1/4 4 Question 6.6. Solve. 3|s| = 18 s = –6 only s = 6 only<< s = –6 or s = 6 s = –18 or s = 18

OpenStudy (anonymous):

Part 2 Question 7.7. Solve. 2x + 6 – 6x = 30 –6<<<< –3 3 6 Question 8.8. Solve. 3(x – 2) + 5x = 18 all real numbers 1 2 3<<<< Question 9.9. Solve. 6 – 4x = 4(3 – x) –2 2 all real numbers no solutions<<< Question 10.10. The length of a rectangle is 5 more than 3 times its width. The perimeter of the rectangle is 58 inches. What is the length of the rectangle? 23 in. 29 in. 32 in.<<<< 35 in. Question 11.11. Which equation has no solution? (Points : 4) -4|x|=20 -3|x|=-15 3|x|=0<<<<<< 4|x|=4 Question 12.12. Solve the formula for the variable h. V=Bh/3 h=3V-B h=3V+B h= V/3B h=3V/B

OpenStudy (welshfella):

re check question 4.4

OpenStudy (welshfella):

and 6.6

OpenStudy (welshfella):

others look OK

OpenStudy (anonymous):

4 : 3/2 ? 6: The third one?

OpenStudy (welshfella):

3/2 is correct

OpenStudy (welshfella):

6 or -6 is correct

OpenStudy (anonymous):

ok thank you. I thought i answered that before and was incorrect so i was unsure thanks!

OpenStudy (welshfella):

yes the value in the mod sign can be pos or neg

OpenStudy (welshfella):

cant check any more gotta go